# Super-intelligence or Superstition? Exploring Psychological Factors Underlying Unwarranted Belief in AI Predictions

### v2

@@ 6 Conclusion
Our study empirically investigated the phenomenon of “rational superstition” in AI by providing a side-by-side comparison of people’s perceptions of fictitious predictions based on AI, astrology, and personality. Even without any validation of system performance, people perceived fictitious AI predictions as highly believable, and their perceptions were positively associated with their perception of astrology- and personality-based predictions. Our findings showed that people’s belief in AI predictions was positively associated with paranormal beliefs, positive attitude towards AI, interest in the topic of prediction, and negatively associated with conscientiousness and age. We also explored interactions with other prediction sources and the subscales including perceived validity, reliability, usefulness, and personalization. In our discussion, we highlighted the role of mental models and cognitive biases in shaping people’s perceptions of AI predictions, and the importance of including psychological and contextual factors in studying human-AI interaction.
@@ 7.2 Experiment Protocol
Figure 3: Dashboard of the simulated investment game
@@ 7.2.1 Assessments and simulated investment game
7.2.2 Predictions
After the game, participants were presented with pre-determined, fictitious predictions from three different sources (AI, astrology, personality) in randomized order. Participants were randomly assigned to either the “Positive” prediction group (rational investor, higher returns) or “Negative” prediction group (impulsive investor, lower returns). Examples of predictions received are shown in Table 2. The initial predictions were written by the authors, with variations generated using ChatGPT-4. These variations were randomly presented to enhance variability and simulate a placebo effect of AI-generated outputs. For each prediction, they were asked to respond to evaluate its perceived validity, personalization, reliability, and usefulness. The evaluation questionnaire is provided in Section 7.3.1.
@@ 7.2.3 Post-study questionnaire
Gullibility A self-reported measure of participants’ tendency to be gullible, developed by [91]. Higher score indicates greater gullibility.
@@ 7.3 Measures
This section includes the detailed questionnaires that we used for the study.
@@ 7.3.1 Perceived validity, reliability, usefulness, and personalization
To measure the believability of the predictions, we presented a short questionnaire to assess perceived validity, reliabiliy, usefulness, and personalization of given predictions. These factors were chosen based on existing literature on the different qualities of a statement that contribute to persuasion, user adoption of an information technology, and positive attitudes.
Perceived validity/accuracy of the content and reliability of the source was added according to the Elaboration Likelihood Model (ELM) in sociology, which argues that the logical validity of an argument is considered central to its believability, but people can also engage in peripheral ways of being persuaded, such as the reliability of the source [93]. Moreover, perceived usefulness was included as a metric that contributes to believability of a prediction, inspired by the Technology Acceptance Model (TAM)[94]. According to this model, perceived usefulness, defined as the degree to which a person believes that using a particular system would enhance their job performance, significantly influences users’ attitudes and intentions, which in turn affect actual technology usage. Lastly, perceived personalization has been found to increase attention and positive attitudes toward a statement [95].
@@ 7.3.2 Cognitive Reflection (CRT-2) and Need for Cognition (NCS-6)
The Need for Cognition (NFC) test is a self-report measure of an individual’s tendency to engage in and enjoy effortful thinking [97]. We use the 6-item NFC assessment (NCS-6) developed by [39].
@@ 7.5.3 Mixed effects model
The simulated investment game is publicly accessible at https://prophecy-study-1.replit.app and the source code is available on GitHub (https://github.com/mitmedialab/ai-prophecy).

### v3

@@ 6 Conclusion
Through a study with 238 participants, this research empirically investigates the phenomenon of “rational superstition” in AI. It provides a side-by-side comparison of people’s perceptions of fictitious predictions based on AI, astrology, and personality. The results suggests that even without any validation of system performance, people perceived fictitious AI predictions as highly believable, and their perceptions were positively associated with their perception of astrology- and personality-based predictions. Our findings showed that people’s belief in AI predictions was positively associated with paranormal beliefs, positive attitude towards AI, interest in the topic of prediction, and negatively associated with conscientiousness and age. We also explored interactions with other prediction sources and subscales of believability, such as perceived validity, reliability, usefulness, and personalization. In our discussion, we highlighted the role of mental models and cognitive biases in shaping people’s perceptions of AI predictions, and the importance of including psychological and contextual factors in studying human-AI interaction.
@@ 7.2 Experiment Protocol
Figure 6: Diagram of the study procedure
@@ 7.2.1 Assessments and simulated investment game
Figure 7: Dashboard of the simulated investment game
7.2.2 Predictions and believability evaluation
After the game, participants were presented with pre-determined, fictitious predictions from three different sources (AI, astrology, personality) in randomized order. Participants were randomly assigned to either the “Positive” prediction group (rational investor, higher returns) or “Negative” prediction group (impulsive investor, lower returns). Examples of predictions received are shown in Table 2. The initial predictions were written by the authors, with variations generated using ChatGPT-4. These variations were randomly presented to enhance variability and simulate AI-generated outputs, extending previous research on placebo effects of AI [69, 70, 71, 72]. For each prediction, participants were asked to evaluate its perceived validity, personalization, reliability, and usefulness. The evaluation questionnaire and rationale are provided in Section 7.3.1.
@@ 7.2.3 Post-study questionnaire
Gullibility A self-reported measure of participants’ tendency to be gullible, developed by Teunisse et al. [100]. Higher score indicates greater gullibility.
@@ 7.3 Measures
This section includes the detailed questionnaires that were used in the study.
@@ 7.3.1 Perceived validity, reliability, usefulness, and personalization
To measure the believability of predictions, we employed four subscales: perceived validity, reliability, usefulness, and personalization. These subscales were informed by theoretical and empirical research on persuasion, technology acceptance, and user attitudes. Together, they capture distinct but interrelated dimensions of how users evaluate predictions from various sources.
Perceived validity reflects the degree to which users find predictions logically sound and factually accurate. This construct is rooted in the Elaboration Likelihood Model (ELM), which proposes that when individuals are motivated and able to process information, they engage in the central route of persuasion, where the quality and validity of arguments are critical determinants of attitude change [102].
Perceived reliability measures the extent to which users view the source of the predictions as trustworthy and dependable. The concept of source reliability aligns with theories of source credibility, which identify trustworthiness and expertise as core elements influencing users’ acceptance of information [103]. Within the ELM framework, the reliability of the source operates as a peripheral cue when individuals are not motivated to critically evaluate the content [102].
Perceived usefulness reflects the degree to which users believe that the predictions can enhance their decision-making or outcomes. This dimension is derived from the Technology Acceptance Model (TAM), where perceived usefulness significantly influences attitudes and intentions toward technology adoption [104]. In the context of predictions, if individuals find the information relevant and applicable, they are more likely to deem it believable because they address users’ goals and needs.
Perceived personalization captures the extent to which users perceive the predictions as tailored to their unique context or characteristics. Personalized communication has been shown to enhance believability by increasing attention and positive attitudes toward the message [105]. Personalization also increases persuasion by making users feel understood and valued, and creates a sense of relevance, which can reduce cognitive load and enhance user engagement [106, 107].
@@ 7.3.2 Cognitive Reflection (CRT-2) and Need for Cognition (NCS-6)
The Need for Cognition (NFC) test is a self-report measure of an individual’s tendency to engage in and enjoy effortful thinking [109]. We use the 6-item NFC assessment (NCS-6) developed by Lins de Holanda Coelho et al. [40].
@@ 7.5.3 Mixed effects model
8 Data Availability
The datasets generated during and analysed during the current study are available in the GitHub repository, https://github.com/mitmedialab/ai-superstition.
9 Code Availability
The source code for the simulated investment game and analysis is available in the GitHub repository, https://github.com/mitmedialab/ai-superstition.
@@ Acknowledgments
10 Author Contributions
E.L. and P.P. conceptualized and designed the study. E.L. implemented the experiment, conducted the study, and performed data analysis and interpretation, with contributions from P.P. E.L. drafted the manuscript and P.P. provided critical revisions. J.A. and P.M. provided feedback throughout the process and edited the manuscript. All authors approved the final version of the manuscript.